# A class of pseudorandom sequences From Function Fields

### v1

@@ I Introduction
We now review several approaches to constructing sequences using the framework of function fields. A standard technique relies on the cyclic structure of finite fields. Let $\mathbb{F}_{q}$ be a finite field with $q$ elements, and let $\alpha$ be a primitive element of $\mathbb{F}_{q}$, where $q=p^{e}$ with $e\geq 1$ and $p$ prime. For any polynomial $f(x)\in\mathbb{F}_{q}[x]$, define the sequence $S_{f}$ by
@@ II-A The linear complexity and correlation
II-B Function fields and exponential sums
@@
    $\left|\sum_{P\in A}\omega_{p}^{\mathrm{Tr}\left(f(P)\right)}\right|\leq\left(\#\mathrm{Supp}((f)_{\infty})+\deg(f)\right)q^{1/2}.$
@@
    $\left|\sum_{P\in A}\omega_{p}^{\mathrm{Tr}\left(f(P)\right)}\right|\leq 2\deg(f)q^{1/2}.$
@@
II-C Elliptic Function Fields
@@
III The general construction of $p$-ary sequences
@@
From the preceding results, we immediately obtain a corollary concerning perfect sequences, showing that the construction can actually be perfect when working over rational function fields or cyclic elliptic function fields.
@@
$\left|N_{(\overline{z},\overline{t})}-\frac{n}{p^{r}}\right|\leq(2g(F)-2)r\sqrt{q}+dr\sqrt{q}(m^{*}_{Q}(z)+1)(1-\frac{1}{p}).$
@@
$\begin{split}&\sum_{(d_{1},\cdots,d_{r})\neq(0,\cdots,0)}\left|\sum^{n-1}_{i=0}\omega_{p}^{d_{1}s_{i+t_{1}}+\cdots+d_{r}s_{i+t_{r}}}\right|\\ &\leq\sum^{r}_{k=1}\binom{r}{k}(p-1)^{k}\left(2g(F)-2+\sum^{k}_{i=1}(m^{*}_{\sigma^{-t_{i}}(Q)}(\sigma^{-t_{i}}(z))+1)d\right)\sqrt{q}\\ &=(2g(F)-2)\sqrt{q}\sum_{k=1}^{r}\binom{r}{k}(p-1)^{k}+d\sqrt{q}(m^{*}_{Q}(z)+1)\sum_{k=1}^{r}\binom{r}{k}k(p-1)^{k}\\ &=(2g(F)-2)r\sqrt{q}p^{r}+dr\sqrt{q}(m^{*}_{Q}(z)+1)(p-1)p^{r-1}.\end{split}$
@@
$\left|f_{(\overline{z},\overline{t})}-\frac{1}{p^{r}}\right|\leq\frac{(2g(F)-2)r\sqrt{q}}{n}+\frac{dr\sqrt{q}(m^{*}_{Q}(z)+1)(p-1)}{pn}.$
@@
Nonlinear complexity is likewise a significant measure for sequences over function fields, and related results can be found in [16, 11, 12]. We now examine this measure for the previously introduced sequences, which can be viewed as a generalization of Theorem III.8 in [12].
@@
Combining these bounds yields $n-\ell\leq p^{h-1}\cdot m^{*}_{Q}(z)\cdot d\cdot(1+\mathrm{L}m)$, and thus

### v2

@@ I Introduction
We now review several approaches to constructing pseudorandom sequences using the framework of function fields. A standard technique relies on the cyclic structure of finite fields. Let $\mathbb{F}_{q}$ be a finite field with $q$ elements and let $\alpha$ be a primitive element of $\mathbb{F}_{q}$, where $q=p^{e}$ with $e\geq 1$ and $p$ prime. For any polynomial $f(x)\in\mathbb{F}_{q}[x]$, define the sequence $S_{f}$ by
@@ II-A The linear complexity and correlation
II-B Nonlinear Complexity
We denote the polynomial ring formed by $L$ variables over finite field by $\mathbb{F}_{q}$ $\mathbb{F}_{q}[x_{1},\ldots,x_{\mathrm{L}}]$. For a monomial $x_{1}^{e_{1}}\cdots x_{\mathrm{L}}^{e_{\mathrm{L}}}$, we denote by $I=(e_{1},\cdots,e_{\mathrm{L}})\in\mathbb{Z}^{L}_{\geq 0}$, with its total degree calculated as $\mathrm{wt}(I)=\sum_{i=1}^{L}e_{i}$. Any element in $\mathbb{F}_{q}[x_{1},...,x_{\mathrm{L}}]$ with a total degree not exceeding $m$ can be expressed as $\sum_{\mathrm{wt}(I)\leq m}a_{I}x_{1}^{e_{1}}\cdots x_{\mathrm{L}}^{e_{\mathrm{L}}},$ where all coefficients $a_{I}$ are elements of $\mathbb{F}_{q}$.
For $\mathbf{s}=(s_{0},s_{1},\cdots,s_{N-1})\neq\mathbf{0}$ defined over $\mathbb{F}_{q}$, the multivariate polynomial $\Phi(x_{1},\cdots,x_{\mathrm{L}})$ generates $\mathbf{s}$ if $s_{i+\mathrm{L}}=\Phi(s_{i},s_{i+1},\cdots,s_{i+\mathrm{L-1}})$ for $0\leq i\leq N-1-\mathrm{L}$. Fix an integer $m$, we denote the $m$th-order nonlinearity by $\mathrm{NL}_{m}(\mathbf{s})$ which associates a minimal positive integer $\mathrm{L}$ where a multivariate polynomial $\Phi(x_{1},\cdots,x_{\mathrm{L}})$ with total degree $\leq m$ can be found to generate the sequence $\mathbf{s}$. Particularly, if each component of $\mathbf{s}$ is zero , $\mathrm{NL}_{m}(\mathbf{s})$ is defined as zero. For $1\leq n\leq N$, $\pi_{n}(\mathbf{s})$ is used to denote the truncated sequence $(s_{0},s_{1},\cdots,s_{n-1})$ which is derived from $\mathbf{s}$ and the $m$th-order nonlinear complexity of this sequence is denoted by $\mathrm{NL}_{m}(\mathbf{s},n)$.
II-C Function fields and exponential sums
@@
    $\left|\sum_{P\in A}\omega_{p}^{\mathrm{Tr}\left(f(P)\right)}\right|\leq\left(\#\mathrm{Supp}((f)_{\infty})+\deg((f)_{\infty})\right)q^{1/2}.$
@@
    $\left|\sum_{P\in A}\omega_{p}^{\mathrm{Tr}\left(f(P)\right)}\right|\leq 2\deg((f)_{\infty})q^{1/2}.$
@@
II-D Elliptic Function Fields
@@
III The general construction of the pseudorandom sequences
@@
From the preceding results, we immediately obtain a corollary concerning perfect sequences, showing that the construction can actually be perfect when working over rational function fields or cyclic elliptic function fields and we shall list them in the last section.
@@
$\left|N_{(\overline{z},\overline{t})}-\frac{n}{p^{r}}\right|\leq(2g(F)-2)\sqrt{q}+dr\sqrt{q}(m^{*}_{Q}(z)+1)(1-\frac{1}{p}).$
@@
$\begin{split}&\sum_{(d_{1},\cdots,d_{r})\neq(0,\cdots,0)}\left|\sum^{n-1}_{i=0}\omega_{p}^{d_{1}s_{i+t_{1}}+\cdots+d_{r}s_{i+t_{r}}}\right|\\ &\leq\sum^{r}_{k=1}\binom{r}{k}(p-1)^{k}\left(2g(F)-2+\sum^{k}_{i=1}(m^{*}_{\sigma^{-t_{i}}(Q)}(\sigma^{-t_{i}}(z))+1)d\right)\sqrt{q}\\ &=(2g(F)-2)\sqrt{q}\sum_{k=1}^{r}\binom{r}{k}(p-1)^{k}+d\sqrt{q}(m^{*}_{Q}(z)+1)\sum_{k=1}^{r}\binom{r}{k}k(p-1)^{k}\\ &=(2g(F)-2)\sqrt{q}p^{r}+dr\sqrt{q}(m^{*}_{Q}(z)+1)(p-1)p^{r-1}.\end{split}$
@@
$\left|f_{(\overline{z},\overline{t})}-\frac{1}{p^{r}}\right|\leq\frac{(2g(F)-2)\sqrt{q}}{n}+\frac{dr\sqrt{q}(m^{*}_{Q}(z)+1)(p-1)}{pn}.$
@@
Nonlinear complexity is likewise a significant measure for sequences over function fields, and related results can be found in [16, 11, 12].
@@
Combining these bounds yields $n-\mathrm{L}\leq p^{h-1}\cdot m^{*}_{Q}(z)\cdot d\cdot(1+\mathrm{L}m)$, and thus